# SciAssess: Benchmarking LLM Proficiency in Scientific Literature Analysis

### v3

@@ Abstract
Recent breakthroughs in Large Language Models (LLMs) have revolutionized natural language understanding and generation, sparking significant interest in applying them to scientific literature analysis. However, existing benchmarks fail to adequately evaluate the proficiency of LLMs in this domain, particularly in scenarios requiring higher-level abilities beyond mere memorization and the handling of multimodal data. In response to this gap, we introduce SciAssess, a benchmark specifically designed for the comprehensive evaluation of LLMs in scientific literature analysis. SciAssess aims to thoroughly assess the efficacy of LLMs by focusing on their capabilities in Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3). It encompasses a variety of tasks drawn from diverse scientific fields, including fundamental science, alloy materials, biomedicine, drug discovery, and organic materials. To ensure the reliability of SciAssess, rigorous quality control measures have been implemented, ensuring accuracy, anonymization, and compliance with copyright standards. SciAssess evaluates 11 LLMs, including GPT, Claude, and Gemini, highlighting their strengths and areas for improvement. This evaluation supports the ongoing development of LLM applications in the analysis of scientific literature. SciAssess and its resources are available at https://sci-assess.github.io/.
@@ 1 Introduction
Figure 1: Overview of SciAssess. It spans over 5 sub-domains and encompasses 29 tasks.
Recent advances in Large Language Models (LLMs), such as GPT-4 [1], Gemini [2], and Llama [3], have attracted considerable attention due to their profound capabilities in natural language understanding and generation [4]. Evaluating these models is crucial for exploring their capability boundaries and limitations, thereby driving technological advancements. In response, a variety of benchmarks tailored for LLMs have been proposed for extensive evaluation, covering a wide range of skills [5, 6] and diverse tasks [7, 8].
@@
Despite LLMs not yet fully replacing scientific researchers in generating creative discoveries, they have demonstrated substantial potential in enhancing researchers’ efficiency in scientific literature analysis [9]. Specific applications such as automatic literature summarization and knowledge extraction have seen practical deployments, significantly boosting researchers’ productivity and expanding the range of literature that can be effectively utilized [10]. Inspired by Bloom’s Taxonomy [11], we systemize the requirements for scientific literature analysis assistants into three progressive levels: (1) Memorization (L1): Establishing an extensive foundational knowledge base to accurately address common factual questions in various scientific domains; (2) Comprehension (L2): Identifying, extracting, and understanding the core content of provided documents; and (3) Analysis & Reasoning (L3): Integrating extracted information with the existing knowledge base to perform logical reasoning and analysis.
@@
Existing comprehensive LLMs benchmarks such as MMLU [12], CMMLU [13], and Xiezhi [14], include some tasks related to scientific data. However, these sub-tasks have two limitations: (1) the tasks mostly focus on Memorization, neglecting higher-level abilities such as L2 and L3; (2) these tasks lack the evaluation of various multi-modal inputs (e.g., charts, molecular structures, and tables), which are of significant presence in scientific literature.
@@
In light of these existing limitations, we introduce SciAssess (cf. Figure 1) – a new benchmark specifically designed for scientific literature analysis. SciAssess not only broadens the evaluation scope to encompass a wider range of LLM capabilities but also extends beyond text to include the extraction and interpretation of multimodal content. Moreover, meticulous design is essential to creating evaluations that yield deep insights, ensure fairness across different LLMs, and provide high utility for stakeholders interested in selecting and optimizing LLMs. Consequently, the design of SciAssess is founded on three critical considerations:
@@
  • Model Ability. A benchmark must clearly delineate the desired capabilities and model the intrinsic relationships among them, facilitating a diagnostic understanding of how these abilities can be acquired and enhanced. Thus, SciAssess evaluates across three progressive levels (i.e., Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3)) and five modalities (i.e., texts, charts, chemical reactions, molecular structures, and tables). Consequently, SciAssess yields nuanced and informative evaluation outcomes, pinpointing specific aspects where the examined models may fall short.
@@
  • Scope & Task. Benchmarks should encompass a broad array of scientific domains to ensure comprehensive representation of disciplines. Within each domain, the selected tasks must authentically represent the typical challenges and scenarios characteristic of that field. To meet this challenge, SciAssess spans over 5 sub-domains (i.e., fundamental science, alloy materials, biomedicine, drug discovery, and organic materials) and encompasses 29 tasks, each carefully suggested or designed by domain practitioners according to their professional experience.
@@
  • Scale & Quality Control. The scale and quality of the benchmark dataset must be impeccable to serve as a dependable basis for deriving accurate, actionable, and applicable insights. SciAssess contains 14,721 questions in total to ensure adequate scale. Each question is transformed from existing datasets or manually curated by domain experts hired by us. Subsequently, rigorous expert cross-validation is performed to ensure correctness and reliability. Additionally, meticulous screening is conducted to remove or anonymize sensitive information, thus protecting privacy and security. These measures guarantee compliance with copyright laws, upholding SciAssess’s legal and ethical integrity.
@@
Overall, SciAssess aims to reveal the performance of LLMs as a scientific literature analysis assistant, thereby identifying both their strength and weaknesses. Following this, our objective is to promote the evolution of LLMs, ensuring they are not only better skilled at navigating scientific literature but also more capable of aiding the progress of research in diverse scientific fields. The insights gained from SciAssess could hopefully catalyze further enhancing the capabilities of LLMs in scientific literature analysis, ultimately contributing to the acceleration of scientific discovery and innovation.
@@ 2 Benchmark Dataset
Table 1: Statistics of the SciAssess benchmark. SciAssess comprises 14,721 questions across 29 tasks in five sub-domains. Tasks are categorized into three ability levels: Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3). The context column denotes whether the task is based on given context from academic papers. Tasks that are gray are transformed from existing public datasets, while others are original, i.e., curated by domain experts hired by us.
Domain Task Ability # Questions Context Question Type Metric Modality
Fundamental Science MMLU (science) L1 2,091 Multiple Choice Accuracy Text only
CMMLU (science) L1 1,700 Multiple Choice Accuracy Text only
Xiezhi-Ch (science) L1 2,882 Multiple Choice Accuracy Text only
Xiezhi-En (science) L1 2,882 Multiple Choice Accuracy Text only
Alloy Materials Alloy Chart QA L2 15 Multiple Choice Accuracy Chart
Composition Extraction L2 244 Table Extraction Table Accuracy Table
Temperature Extraction L2 207 Multiple Choice Accuracy Text only
@@
Treatment Sequence L3 102 True/False Accuracy Text only
Biomedicine Biology Chart QA L2 99 Multiple Choice Accuracy Chart
Chemical Entities Recognition L2 997 Text Extraction Recall Text only
Disease Entities Recognition L2 997 Text Extraction Recall Text only
Compound Disease Recognition L3 997 Text Extraction Recall Text only
Gene Disease Function L3 236 Text Extraction Recall Text only
Gene Disease Regulation L3 240 Text Extraction Recall Text only
Drug Discovery Affinity Extraction L2 40 Table Extraction Table Accuracy Mol., Table
@@
Tag to Molecule L2 50 Molecule Generation Mol. Similarity Mol.
Markush to Molecule L3 37 Molecule Generation Mol. Similarity Mol.
@@
Reaction QA L3 95 Multiple Choice Accuracy Reaction
Drug Target Identification L3 40 Text Extraction Recall Text only
Organic Materials Electrolyte Table QA L2 100 Multiple Choice Accuracy Table
OLED Property Extraction L2 13 Table Extraction Table Accuracy Mol.,Table
Polymer Chart QA L2 15 Multiple Choice Accuracy Chart
Polymer Composition QA L2 109 Multiple Choice Accuracy Text only
Polymer Property Extraction L2 109 Table Extraction Table Accuracy Table
Solubility Extraction L2 100 Table Extraction Table Accuracy Table
Reaction Mechanism QA L3 22 Multiple Choice Accuracy Reaction
@@ 2.1 Ability Assessment Framework
Guided by the widely accepted cognitive learning processes outlined in Bloom’s Taxonomy [11], we propose that the evaluation of LLMs in scientific literature analysis should be classified into three core levels:
  • Memorization (L1) refers to the model’s extensive knowledge base, which allows it to accurately answer common factual questions in science autonomously.
  • Comprehension (L2) is the ability to precisely identify and extract key information and facts within a given text, and to comprehend them.
  • Analysis & Reasoning (L3) demonstrate the model’s advanced capability to amalgamate extracted information with its existing knowledge base for logical reasoning and analysis, leading to well-founded conclusions or predictions.
Inspecting existing LLM benchmarks [12, 13, 14] in science field through three-level ability assessment framework, we find that they mostly focus on Memorization (L1) – the foundational knowledge base for scientific facts – while overlooking the higher-level abilities of Comprehension (L2) and Analysis & Reasoning (L3).
@@ 2.2 Scope & Task
After categorizing the ability of of LLMs into three levels, we proceed to introduce how we choose the tasks in SciAssess. As mentioned above, Memorization (L1), being the extensive foundation for other higher-level abilities, should encompass as large a knowledge base as possible. Therefore, SciAssess initially includes factual questions covering fundamental science subjects, such as chemistry, biology, etc.. For the evaluation of Comprehension (L2) and Analysis & Reasoning (L3), we focus on vertical domains such as materials and biomedicine. The reason is that solving tasks in these domains require finer-grained abilities, such as understanding tables and molecular structures. For instance, crucial composition information in alloy material literature is often found in tables, whereas key information extraction in drug discovery necessitates the accurate recognition of molecular structures. Thus, we further include four vertical domains: alloy materials, biomedicine, drug discovery, and organic materials, as shown in Figure 1. This categorization ensures that the evaluation framework captures the unique challenges and requirements of each specific field.
@@
SciAssess, as presented in Table 1, comprises 14,721 questions across 29 tasks in five practical scientific disciplines, encompassing three ability levels: Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3). This ensures a comprehensive evaluation of LLMs in scientific literature analysis. Of these tasks, 19 out of 29 are based on contexts from academic papers, specifically designed to assess the models’ ability to analyze scientific literature. We show the token lengths (in GPT-4 tokenizer) of questions and answers for each task in Figure 3. SciAssess also includes five types of questions (i.e., True/False questions, Multiple-Choice questions, Table Extraction, Text Extraction, and Molecule Generation) with four metrics (i.e., accuracy, table accuracy, recall, molecule similarity). For an in-depth understanding of these question types and metrics, including detailed descriptions and concrete examples, please refer to Appendix A. We also provide prompt template and specific prompt for each task in Appendix B and C, respectively.
2.2.1 Fundamental Science
Aiming to evaluate foundational knowledge – the basis of literature analysis, fundamental science should encompass as large a knowledge base as possible. Several widely-adopted benchmarks could be leveraged to evaluate LLMs’ awareness of commonsense facts across various domains. Specifically, MMLU [12] consists of 57 tasks on various subjects, each containing multiple-choice questions that test domain-specific facts. CMMLU [13] is a comprehensive Chinese benchmark that covers a wide range of 67 subjects. Xiezhi [14] is a Chinese-English bilingual benchmark with 249,587 multiple-choice questions spanning 516 diverse disciplines. To evaluate LLMs’ capabilities in fundamental science, we filter out the STEM subsets, excluding mathematics, from these benchmarks and adapt them into a unified multi-choice template. This ensures a robust assessment of the models’ foundational knowledge, which is crucial for their performance in scientific literature analysis.
2.2.2 Alloy Materials
Alloy materials are composed of two or more metal elements mixed in specific proportions to exhibit metal-like characteristics. These materials are extensively used in aerospace, automobile manufacturing, construction, and other industries. By adjusting their composition and preparation processes, alloy materials can be tailored to meet specific properties and requirements [15]. Thus, accurately extracting alloy composition and process parameters from literature is crucial for effective alloy design.
In this domain, we evaluate the capabilities of LLMs in extracting key information for alloy design. We have devised a comprehensive set of tasks related to literature research, which include extracting alloy composition, extracting process values, determining process sequences, and differentiating between samples. The standard answers for all tasks have been manually extracted from literature across various journals and subsequently verified by an independent expert.
Alloy Chart QA The processing steps and properties of alloy materials are often presented in charts, such as those comparing the performance of multiple alloys or illustrating how elongation changes with composition. Therefore, extracting information from these charts and integrating it with textual information is crucial. To further evaluate the retrieval capability of models regarding alloy chart information, we have designed multiple-choice questions involving alloy composition, processing techniques, and properties.
Composition Extraction Extracting alloy composition information from an article’s text or tables and unifying it into a structured format helps researchers utilize historical data more effectively and provides valuable guidance for subsequent designs. This comprehensive task evaluates LLMs’ ability to extract alloy compositions (including all element contents) from text and tables. Typically, alloy element content is found in two cases: (1) the element content is stored in a table, and (2) the element content is implicitly indicated by the alloy name, such as ‘Fe30Co20Ni50’, which represents an atomic ratio of 30% Fe, 20% Co, and 50% Ni. The objective of this task is to comprehensively extract this information and organize it into a digestible table. The metric is to calculate the matching score between the standard answer table and the extraction result table. This task showcases the LLM’s comprehension ability to integrate, extract, and structure multi-modal information [16].
Temperature Extraction The properties of an alloy are determined by its composition and the processes it undergoes, including processing and heat treatment. Therefore, extracting heat treatment values is critical. This task aims to determine the maximum temperature value for the heat treatment of the alloy. To ensure easy statistical analysis, questions are designed as multiple-choice.
Sample Differentiation Alloys with the same composition but treated by different processes are considered different samples because they exhibit different properties. Therefore, distinguishing between different samples and understanding the differences in their processes is essential. This multiple-choice question task is designed to comprehensively judge the number of different alloy samples proposed or studied by the authors. It assesses the LLMs’ analysis and reasoning abilities regarding alloy distinctions from text.
Treatment Sequence Each alloy treatment process has a clear sequence requirement, so it is necessary to ensure that the extracted heat treatment process sequence is consistent with the experimental sequence. For example, after solution treatment, a sample is further aged to ensure the release of internal stresses. This task aims to objectively analyze and evaluate the sequential relationship between two heat treatments and provide True/False answers. Additionally, if a specific heat treatment name does not exist in the paper, it should be considered False. This task assesses the LLM’s comprehension ability to judge treatment order from the text.
2.2.3 Biomedicine

### v5

@@ Abstract
Recent breakthroughs in Large Language Models (LLMs) have revolutionized scientific literature analysis. However, existing benchmarks fail to adequately evaluate the proficiency of LLMs in this domain, particularly in scenarios requiring higher-level abilities beyond mere memorization and the handling of multimodal data. In response to this gap, we introduce SciAssess, a benchmark specifically designed for the comprehensive evaluation of LLMs in scientific literature analysis. It aims to thoroughly assess the efficacy of LLMs by evaluating their capabilities in Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3). It encompasses a variety of tasks drawn from diverse scientific fields, including biology, chemistry, material, and medicine. To ensure the reliability of SciAssess, rigorous quality control measures have been implemented, ensuring accuracy, anonymization, and compliance with copyright standards. SciAssess evaluates 11 LLMs, highlighting their strengths and areas for improvement. We hope this evaluation supports the ongoing development of LLM applications in scientific literature analysis. SciAssess and its resources are available at https://github.com/sci-assess/SciAssess.
SciAssess: Benchmarking LLM Proficiency in Scientific Literature Analysis
Hengxing Cai1thanks: Equal Contribution, Xiaochen Cai [^1] footnotemark: 1, Junhan Chang [^1] footnotemark: 1, Sihang Li [^1] footnotemark: 1, Lin Yao1, Changxin Wang1, Zhifeng Gao1, Hongshuai Wang1, Yongge Li1, Mujie Lin1, Shuwen Yang1, Jiankun Wang1, Mingjun Xu1, Jin Huang1, Xi Fang1, Jiaxi Zhuang1, Yuqi Yin1, Yaqi Li1, Changhong Chen1, Zheng Cheng2, Zifeng Zhao2, Linfeng Zhang1,2 and Guolin Ke1 1DP Technology  2AI for Science Institute, Beijing
@@ 1 Introduction
Figure 1: Overview of SciAssess. It spans over 4 sub-domains and encompasses 27 tasks.
Recent advances in Large Language Models (LLMs), such as GPT OpenAI (2023); Brown (2020), Gemini Google (2023), and Llama Touvron et al. (2023), have attracted considerable attention due to their profound capabilities in natural language understanding and generation Bubeck et al. (2023). Evaluating these models is crucial for exploring their capability boundaries and limitations, thereby driving technological advancements. In response, a variety of benchmarks tailored for LLMs have been proposed for extensive evaluation, covering a wide range of skills Zhong et al. (2023); Huang et al. (2023) and diverse tasks Srivastava et al. (2022); Suzgun et al. (2023a).
@@
Despite LLMs not yet fully replacing scientific researchers in generating creative discoveries, they have demonstrated substantial potential in enhancing researchers’ efficiency in scientific literature analysis AI4Science and Quantum (2023). Specific applications such as automatic literature summarization and knowledge extraction have seen practical deployments, significantly boosting researchers’ productivity and expanding the range of literature that can be effectively utilized Zheng et al. (2023). Inspired by Bloom’s Taxonomy Krathwohl (2002), we systemize the requirements for scientific literature analysis assistants into three progressive levels: (1) Memorization (L1): Establishing an extensive foundational knowledge base to accurately address common factual questions in various scientific domains; (2) Comprehension (L2): Identifying, extracting, and understanding the core content of provided documents; and (3) Analysis & Reasoning (L3): Integrating extracted information with the existing knowledge base to perform logical reasoning and analysis.
@@
Existing comprehensive LLMs benchmarks, such as MMLU-Pro Hendrycks et al. (2021); Wang et al. (2024), include some tasks related to scientific data. However, these sub-tasks have two limitations: (1) they mostly focus on Memorization, neglecting higher-level abilities such as L2 and L3; (2) these tasks lack the evaluation of various multi-modal inputs (e.g., charts, molecular structures, and tables), which are crucial in scientific literature.
@@
In light of these existing limitations, we introduce SciAssess (cf. Figure 1) – a benchmark specifically designed for scientific literature analysis. SciAssess not only broadens the evaluation scope to encompass a wider range of LLM capabilities but also extends beyond text to include the extraction and interpretation of multimodal contents. Moreover, meticulous design is essential to creating evaluations that yield deep insights, ensure fairness across different LLMs. Consequently, SciAssess is founded on three critical considerations:
@@
Model Ability. A benchmark must clearly delineate the desired capabilities and model the intrinsic relationships among them, facilitating a diagnostic understanding. Thus, SciAssess evaluates across three progressive levels (i.e., Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3)) and five modalities (i.e., texts, charts, chemical reactions, molecular structures, and tables). Consequently, SciAssess yields nuanced and informative evaluation outcomes, pinpointing specific aspects where the examined models may fall short.
@@
Scope & Task. Benchmarks should encompass a broad array of scientific domains to ensure comprehensiveness. Within each domain, the selected tasks must authentically represent the typical challenges and scenarios characteristic of that field. Consequently, SciAssess spans over 4 sub-domains (i.e., biology, chemistry, material, and medicine) and encompasses 27 tasks, each carefully suggested or designed by domain experts according to their professional experience.
@@
Scale & Quality Control. The scale and quality of the benchmark must be impeccable to serve as a dependable basis for deriving accurate, actionable, and applicable insights. SciAssess contains 6,938 questions in total to ensure adequate scale. Each question is transformed from existing datasets or manually curated by domain experts hired by us  [^1] All data collection, annotation, and quality control tasks were carried out by the authors (who are also employees of the company) as part of their job responsibilities, and therefore, they were not provided with any additional compensation.. Subsequently, expert cross-validation is performed to ensure correctness and reliability.
@@
Overall, SciAssess aims to reveal the performance of LLMs as a scientific literature analysis assistant, thereby identifying their strength and weaknesses. The insights gained from SciAssess could hopefully catalyze further enhancing the capabilities of LLMs in scientific literature analysis, ultimately contributing to the acceleration of scientific discovery and innovation.
@@ 2 Benchmark Dataset
Domain Task Ability # Questions Question Type Metric Modality
Biology MMLU-Pro-Biology L1 717 Multiple Choice Accuracy Text only
Biology Chart QA L2 199 Multiple Choice Accuracy Chart
Chemical Entities Recognition L2 500 Text Extraction F1-score Text only
Compound Disease Recognition L2 775 Text Extraction F1-score Text only
Disease Entities Recognition L2 500 Text Extraction F1-score Text only
Gene Disease Function L2 24 Text Extraction F1-score Text only
Chemistry MMLU-Pro-Chemistry L1 1,132 Multiple Choice Accuracy Text only
Electrolyte Table QA L2 200 Multiple Choice Accuracy Table
OLED Property Extraction L2 13 Table Extraction Recall Mol., Table
Polymer Chart QA L2 15 Multiple Choice Accuracy Chart
Polymer Composition QA L2 209 Multiple Choice Accuracy Text only
Polymer Property Extraction L2 109 Table Extraction Recall Table
Solubility Extraction L2 100 Table Extraction Recall Table
Reactant QA L3 195 Multiple Choice Accuracy Reaction
Reaction Mechanism QA L3 22 Multiple Choice Accuracy Reaction
Material Material QA L1 263 Multiple Choice Accuracy Text only
Alloy Chart QA L2 15 Multiple Choice Accuracy Chart
Composition Extraction L2 244 Table Extraction Recall Table
Temperature QA L2 207 Multiple Choice Accuracy Text only
@@
Treatment Sequence L3 202 True/False Accuracy Text only
Medicine MMLU-Pro-Health L1 818 Multiple Choice Accuracy Text only
Affinity Extraction L2 40 Table Extraction Recall Mol., Table
@@
Tag to Molecule L2 50 Mol. Generation Mol. Similarity Mol.
Markush to Molecule L3 37 Mol. Generation Mol. Similarity Mol.
@@
Table 1: Statistics of the SciAssess. It comprises 6,888 questions across 27 tasks in five sub-domains. Tasks are categorized into three ability levels: Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3). Tasks that are gray are transformed from existing datasets, while others are curated by domain experts hired by us.
@@ 2.1 Ability Assessment Framework
Guided by the widely accepted cognitive learning processes outlined in Bloom’s Taxonomy Krathwohl (2002), we propose that the evaluation of LLMs in scientific literature analysis should be classified into three core levels:
Memorization (L1) refers to the model’s extensive knowledge base, which allows it to accurately answer common factual questions in science autonomously. Comprehension (L2) is the ability to precisely identify and extract key information and facts within a given text, and to comprehend them. Analysis & Reasoning (L3) demonstrate the model’s advanced capability to amalgamate extracted information with its existing knowledge base for logical reasoning and analysis, leading to well-founded conclusions or predictions.
Inspecting existing LLM benchmarks in science field (See Section 4) through three-level ability assessment framework, we find that they mostly focus on Memorization (L1) – the foundational knowledge base for scientific facts – while overlooking the higher-level abilities of Comprehension (L2) and Analysis & Reasoning (L3).
@@ 2.2 Scope & Task
After categorizing the ability of of LLMs into three levels, we proceed to introduce how we choose the tasks in SciAssess. First, we include four vertical domains: biology, chemistry, material, and medicine, as shown in Figure 1. This categorization ensures that SciAssess captures the unique challenges and requirements of each specific field. Then, as mentioned above, Memorization (L1), being the extensive foundation for other higher-level abilities, should encompass as large a knowledge base as possible. Thus, SciAssess includes factual questions in MMLU-Pro Wang et al. (2024) and MaScQA Zaki et al. (2023), covering fundamental knowledge in each field. For the evaluation of Comprehension (L2) and Analysis & Reasoning (L3), we identify realistic demands by consulting domain experts and curate corresponding tasks. The reason is that solving tasks in these domains require finer-grained abilities, such as understanding tables and molecular structures. For instance, crucial composition information in material science literature is often found in tables, whereas key information extraction in drug discovery necessitates the accurate recognition of molecular structures.
@@
SciAssess, as presented in Table 1, comprises 6,888 questions across 27 tasks in five scientific domain, encompassing three ability levels: Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3). Of these tasks, 7 out of 27 are transformed from existing public datasets (gray tasks in Table 1) , and the other 21 tasks curated by us are based on contents from academic papers, specifically designed to assess the ability to analyze scientific literature. We show the token lengths (GPT-4 tokenizer) of questions and answers for each task in Figure 3. SciAssess also includes five types of questions (i.e., true/false questions, multiple-choice questions, table extraction, text extraction, and molecule generation) with four metrics (i.e., accuracy, recall, F1-score, and molecule similarity). For detailed descriptions and concrete examples, please refer to Appendix A. We also provide general prompt template and specific prompt for each task in Appendix B and C, respectively.
2.2.1 Biology